# Theory for the spectral splitting exponent of exceptional points

### v1

@@ I Introduction
Recently, non-Hermitian Hamiltonians have garnered considerable attention [1, 2] as an effective framework for describing open quantum and classical systems. In such systems, the effective Hamiltonian becomes non-Hermitian due to gain and loss [3, 4]. Notably, the introduction of non-Hermiticity gives rise to numerous phenomena without Hermitian counterparts, such as the emergence of exceptional points (EPs) [5, 6], the non-Bloch band theory [7, 8, 9, 10, 11, 12, 13, 14], non-Hermitian skin effects [7, 8, 15, 16, 17, 18, 12, 19, 14, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], novel non-Hermitian topological phases [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 35, 48, 36, 37], edge bursts [49, 50, 51], non-Bloch $\mathcal{PT}$ symmetry breaking [52, 53], and non-Hermitian invisibility [54, 55].
@@
EPs correspond to degeneracies in non-Hermitian energy spectra [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76]. These spectral singularities give rise to a plethora of exotic physical phenomena, such as unidirectional invisibility [77, 78], topological energy transfer [79, 80] and laser mode selectivity [81, 82]. Due to the rich physics they enable, EPs have been realized experimentally across numerous platforms, including photonic crystals [83, 84, 85, 86], exciton-polariton systems [87], cavities [88, 89], metamaterials [90] and synthetic systems [91].
@@
A salient property of EPs is their sensitive spectral response to perturbations, which is rooted in their singular wave-function structure. Unlike Hermitian degeneracies, where eigenstates remain orthogonal, EPs are characterized by the simultaneous coalescence of both eigenvalues and their corresponding eigenstates. This fundamental distinction leads to significantly different responses to perturbations. Specifically, under a perturbation of strength $\epsilon$, the spectral splitting for a Hermitian degeneracy scales at most linearly with $\epsilon$ ($\Delta E\sim\epsilon$). In stark contrast, the splitting for an $N$th-order EP (where $N$ eigenstates coalesce) can scale as $\Delta E\sim\epsilon^{1/N}$ [92, 93, 94]. The ratio of these responses, $\epsilon^{1/N}/\epsilon=\epsilon^{(1-N)/N}$, diverges in the limit $\epsilon\rightarrow 0$. This indicates that the response of an EP to a perturbation can be much stronger than that of any Hermitian degeneracy, a property that makes EPs exceptional candidates for enhancing sensitivity in sensing applications [95, 96, 93, 94, 97, 98, 99, 100, 101, 102].
@@
While the strongest response, $\Delta E\sim\epsilon^{1/N}$, should in principle yield optimal sensor performance, this scaling is not guaranteed for an arbitrary perturbation [103, 104, 105]. For an $N$th-order EP, the exponent $\alpha$ characterizing the leading-order splitting (defined by $\Delta E\sim\epsilon^{\alpha}$) can range from $\alpha=1$ to $\alpha=1/N$. Specially, the exponent takes the general form $\alpha=1/k$, where $k$ is an integer satisfying $1\leq k\leq N$. The integer $k$ corresponds to the number of eigenstates that escape from the coalescence. Although the exponent $\alpha$ can always be determined in principle by solving the characteristic equation, this approach becomes analytically intractable for large $N$. Furthermore, numerical solutions, while feasible, often fail to provide physical insight into the relationship between the perturbation’s form and the resulting exponent. Therefore, an analytical theory capable of predicting $\alpha$ without solving for the full spectrum would be highly valuable, as it would enable the design of perturbations that yield a desired splitting scaling.
@@
In a previous work, we explored the spectral splitting of infernal points—a class of EP whose order scales with system size—under perturbations [69]. Because the Hamiltonian at these points takes a Jordan form, we were able to analytically derive the condition for achieving the strongest possible response ($\alpha=1/N$). Intriguingly, the condition exhibits a simple relationship to the perturbation term’s positions in the matrix representation. Given that the Jordan block structure is a universal feature of Hamiltonians at EPs, we here extend this theory to general EPs of arbitrary order. We demonstrate that the splitting exponent $\alpha$ for a generic response can be determined analytically when the perturbation matrix satisfies specific algebraic conditions. This establishes a direct connection between the perturbation’s matrix structure and the exponent $\alpha$, with our previous result emerging as a special case [69].
@@ II General framework
We consider a general $N\times N$ non-Hermitian Hamiltonian $H_{\rm N}$ which supports an $N$th-order EP for appropriate parameters. Without loss of generality, we assume the EP occurs at energy $E_{0}$. At this degeneracy, $H_{\rm N}$ possesses a single eigenstate, $|u_{0}\rangle$, satisfying $H_{\rm N}|u_{0}\rangle=E_{0}|u_{0}\rangle$. The complete $N$-dimensional Hilbert space is spanned by the Jordan chain, comprising this eigenstate and its N1N-1 associated vectors: $|u_{0}\rangle,|u_{1}\rangle,\cdots,|u_{N-1}\rangle$}. Correspondingly, the system has a single left eigenstate, $|v_{0}\rangle$, satisfying $H_{\rm N}^{\dagger}|v_{0}\rangle=E_{0}^{*}|v_{0}\rangle$, along with N1N-1 left associated vectors, $|v_{1}\rangle,\cdots,|v_{N-1}\rangle$. Owing to the Hamiltonian’s defectiveness, the standard biorthogonality condition $\langle v_{i}|u_{j}\rangle=\delta_{ij}$ breaks down. Instead, the vectors obey the Jordan chain structure [106], characterized by the relation $\langle v_{i}|u_{j}\rangle=\delta_{i+j,N-1}$ (A brief review is provided in Appendix A).
@@ III.2 Case 2
where $m=1,2,\cdots,N-p-q$. This result indicates that the spectral splitting scales as $\Delta E\sim\epsilon^{1/(N-p-q)}$. Given that $0\leqslant p+q<N$, a single-element perturbation can produce any splitting exponent of the form $\alpha=1/k$ for integers $k$ satisfying $1\leqslant k\leqslant N$. In particular, the choice $p=q=0$ yields the strongest possible scaling, $\Delta E\sim\epsilon^{1/N}$, which recovers the result from our previous work [69].
@@ III.4 Case 4
The consistent result $\alpha=1/(N-p-q)$ across the last three cases leads us to conjecture that the splitting exponent is a function only of the index sum $p+q$. In other words, $\alpha$ is governed exclusively by the locations of the nonzero elements of the perturbation matrix.
@@ IV Application in a concrete model
While many models have been proposed to realize EPs, we select a non-Hermitian supersymmetric (SUSY) array model [92] to illustrate our theoretical framework. This model is ideally suited for this purpose, as it allows for the construction of EPs of any order. Specially, the Hamiltonian for an $N$-site non-Hermitian SUSY array is
@@
Figure 1: Absolute eigenenergy splitting $|\Delta E|$ for the 4-site non-Hermitian SUSY array under five different perturbations $H_{p1}$ to $H_{p5}$. Dots represent numerical results, while solid curves show the analytical predictions $|\Delta E|=\epsilon^{1/4}$, $\epsilon^{1/3}$, $\epsilon^{1/2}$, $\epsilon$, and 0, respectively. Parameters are $\omega_{0}=0$, $\gamma=1$, and $J=1$.
@@ Appendix A A brief review of the Jordan chain structure
Consider an $N$th-order EP at energy $E_{0}$, described by an $N\times N$ non-Hermitian Hamiltonian $H_{N}$. This Hamiltonian possesses a unique eigenstate $|u_{0}\rangle$ and, as shown in Ref. [106], N1N-1 associated vectors $\{|u_{1}\rangle,\dots,|u_{N-1}\rangle\}$ that together form a Jordan chain satisfying:

### v2

@@ I Introduction
Recently, non-Hermitian Hamiltonians have garnered considerable attention [RevModPhys.93.015005, ashida2020non] as an effective framework for describing open quantum and classical systems. In such systems, the effective Hamiltonian becomes non-Hermitian due to gain and loss [Rotter_2009, PhysRevLett.70.2273]. Notably, the introduction of non-Hermiticity gives rise to numerous phenomena without Hermitian counterparts, such as the emergence of exceptional points (EPs) [W.D.Heiss_2004, berry2004physics], the non-Bloch band theory [PhysRevLett.121.086803, PhysRevLett.123.066404, PhysRevB.101.195147, PhysRevX.14.021011, HU202551, zhang2022universal, PhysRevLett.133.216401, PhysRevB.107.115412], non-Hermitian skin effects [PhysRevLett.121.086803, PhysRevLett.123.066404, PhysRevLett.124.056802, PhysRevLett.125.126402, PhysRevLett.124.086801, PhysRevLett.125.226402, zhang2022universal, PhysRevLett.125.186802, PhysRevB.107.115412, PhysRevB.109.L081108, PhysRevB.108.L060204, PhysRevB.106.L241112, li2021impurity, PhysRevResearch.5.033058, PhysRevB.108.075122, PhysRevB.108.205423, PhysRevLett.127.256402, PhysRevLett.131.116601, PhysRevLett.127.116801, PhysRevB.111.L140201, li2020critical, PhysRevB.104.165117, PhysRevB.107.155430, li2024observation, PhysRevLett.123.016805, PhysRevB.102.241202, PhysRevB.103.045420, Chen_2024, manna2023inner], novel non-Hermitian topological phases [PhysRevX.8.031079, PhysRevX.9.041015, PhysRevLett.120.146402, PhysRevB.110.L121401, PhysRevB.99.201103, PhysRevLett.116.133903, PhysRevLett.121.136802, PhysRevLett.126.216405, PhysRevLett.122.076801, PhysRevB.102.205118, PhysRevLett.123.016805, PhysRevB.99.081302, PhysRevB.102.241202, PhysRevB.103.045420, PhysRevB.111.085148, PhysRevB.111.094109, PhysRevB.111.195419], edge bursts [PhysRevLett.128.120401, PhysRevB.108.235422, PhysRevLett.133.070801], non-Bloch $\mathcal{PT}$ symmetry breaking [PhysRevLett.132.050402, PhysRevLett.126.230402], and non-Hermitian invisibility [PhysRevA.82.032111, PhysRevB.106.094205].
@@
EPs correspond to degeneracies in non-Hermitian energy spectra [Heiss_2012, ding2022non, PhysRevResearch.2.043045, PhysRevLett.126.086401, PhysRevB.105.075420, PhysRevB.104.L201104, PhysRevB.104.L161116, PhysRevB.104.L161117, PhysRevLett.123.066405, gohsrich2024, PhysRevLett.128.010402, denner2021exceptional, PhysRevB.110.L201104, wang2025, PhysRevResearch.6.043023, PhysRevLett.127.186602, PhysRevLett.118.045701, PhysRevB.99.081102, PhysRevA.98.042114, PhysRevB.99.161115, PhysRevLett.126.010401, PhysRevB.110.L020104, PhysRevLett.134.133801, 18gg-gvzc, PhysRevResearch.4.023130, bid2025fragmentedexceptionalpointsbulk, PhysRevResearch.7.023062], and share intriguing connections with Newton polygons and tropical geometry [jaiswal2023characterizing, banerjee2023tropical]. These spectral singularities give rise to a plethora of exotic physical phenomena, such as unidirectional invisibility [Lin2011unidirectional, Peng2014], topological energy transfer [Doppler2016, Xu2016] and laser mode selectivity [Hossein2014, Feng2014]. Due to the rich physics they enable, EPs have been realized experimentally across numerous platforms, including photonic crystals [ozdemir2019parity, regensburger2012parity, zhou2023observation, kim2016direct], exciton-polariton systems [gao2015observation], cavities [PhysRevX.6.021007, PhysRevLett.103.134101], metamaterials [Park2020] and synthetic systems [yang2023realization].
@@
A salient property of EPs is their sensitive spectral response to perturbations, which is rooted in their singular wave-function structure. Unlike Hermitian degeneracies, where eigenstates remain orthogonal, EPs are characterized by the simultaneous coalescence of both eigenvalues and their corresponding eigenstates. This fundamental distinction leads to significantly different responses to perturbations. Specifically, under a perturbation of strength $\epsilon$, the spectral splitting for a Hermitian degeneracy scales at most linearly with $\epsilon$ ($\Delta E\sim\epsilon$). In stark contrast, the splitting for an $N$th-order EP (where $N$ eigenstates coalesce) can scale as $\Delta E\sim\epsilon^{1/N}$ [PhysRevA.101.033820, chen2017exceptional, hodaei2017enhanced]. The ratio of these responses, $\epsilon^{1/N}/\epsilon=\epsilon^{(1-N)/N}$, diverges in the limit $\epsilon\rightarrow 0$. This indicates that the response of an EP to a perturbation can be much stronger than that of any Hermitian degeneracy, a property that makes EPs exceptional candidates for enhancing sensitivity in sensing applications [PhysRevLett.112.203901, PhysRevA.93.033809, chen2017exceptional, hodaei2017enhanced, PhysRevLett.123.213901, PhysRevLett.122.153902, PhysRevApplied.12.024002, PhysRevLett.125.240506, PhysRevA.101.053846, PhysRevResearch.6.033284]. The ultimate performance in practical devices, however, is limited by noise, which bounds the signal-to-noise ratio [lau2018fundamental, Chen_2019]. Strategies to overcome this limitation remain an active area of research [Kononchuk2022]. A deeper understanding of the spectral splitting behavior can shed new light on these key problems.
@@
While the strongest response, $\Delta E\sim\epsilon^{1/N}$, should in principle yield optimal sensor performance, this scaling is not guaranteed for an arbitrary perturbation [PhysRevLett.127.186601, PhysRevLett.134.153601, wu2021high, PhysRevB.99.241403]. For an $N$th-order EP, the exponent $\alpha$ characterizing the leading-order splitting (defined by $\Delta E\sim\epsilon^{\alpha}$) can range from $\alpha=1$ to $\alpha=1/N$. Specially, the exponent takes the general form $\alpha=1/k$, where $k$ is an integer satisfying $1\leq k\leq N$. The integer $k$ corresponds to the number of eigenstates that escape from the coalescence. Although the exponent $\alpha$ can always be determined in principle by solving the characteristic equation, this approach becomes analytically intractable for large $N$. Furthermore, numerical solutions, while feasible, often fail to provide physical insight into the relationship between the perturbation’s form and the resulting exponent. Therefore, an analytical theory capable of predicting $\alpha$ without solving for the full spectrum would be highly valuable, as it would enable the design of perturbations that yield a desired splitting scaling.
@@
In a previous work, we explored the spectral splitting of infernal points—a class of EP whose order scales with system size—under perturbations [wang2025]. Because the Hamiltonian at these points takes a Jordan form, we were able to analytically derive the condition for achieving the strongest possible response ($\alpha=1/N$). Intriguingly, the condition exhibits a simple relationship to the perturbation term’s positions in the matrix representation. Given that the Jordan block structure is a universal feature of Hamiltonians at EPs, we here extend this theory to general EPs of arbitrary order. We demonstrate that the splitting exponent $\alpha$ for a generic response can be determined analytically when the perturbation matrix satisfies specific algebraic conditions. This establishes a direct connection between the perturbation’s matrix structure and the exponent $\alpha$, with our previous result emerging as a special case [wang2025].
@@ II General framework
We consider a general $N\times N$ non-Hermitian Hamiltonian $H_{\rm N}$ which supports an $N$th-order EP for appropriate parameters. Without loss of generality, we assume the EP occurs at energy $E_{0}$. At this degeneracy, $H_{\rm N}$ possesses a single eigenstate, $|u_{0}\rangle$, satisfying $H_{\rm N}|u_{0}\rangle=E_{0}|u_{0}\rangle$. The complete $N$-dimensional Hilbert space is spanned by the Jordan chain, comprising this eigenstate and its N1N-1 associated vectors: $|u_{0}\rangle,|u_{1}\rangle,\cdots,|u_{N-1}\rangle$}. Correspondingly, the system has a single left eigenstate, $|v_{0}\rangle$, satisfying $H_{\rm N}^{\dagger}|v_{0}\rangle=E_{0}^{*}|v_{0}\rangle$, along with N1N-1 left associated vectors, $|v_{1}\rangle,\cdots,|v_{N-1}\rangle$. Owing to the Hamiltonian’s defectiveness, the standard biorthogonality condition $\langle v_{i}|u_{j}\rangle=\delta_{ij}$ breaks down. Instead, the vectors obey the Jordan chain structure [seyranian2003multiparameter], characterized by the relation $\langle v_{i}|u_{j}\rangle=\delta_{i+j,N-1}$ (A brief review is provided in Appendix A).
@@ III.2 Case 2
where $m=1,2,\cdots,N-p-q$. This result indicates that the spectral splitting scales as $\Delta E\sim\epsilon^{1/(N-p-q)}$. Given that $0\leqslant p+q<N$, a single-element perturbation can produce any splitting exponent of the form $\alpha=1/k$ for integers $k$ satisfying $1\leqslant k\leqslant N$. In particular, the choice $p=q=0$ yields the strongest possible scaling, $\Delta E\sim\epsilon^{1/N}$, which recovers the result from our previous work [wang2025].
@@ III.4 Case 4
The consistent result $\alpha=1/(N-p-q)$ across the last three cases leads us to conjecture that the splitting exponent is a function only of the index sum $p+q$, i.e., $j$. When the perturbation matrix contains terms with different values of $j$, the minimum value of $j$ governs the splitting exponent in the extremely weak-perturbation regime. In other words, $\alpha$ is governed exclusively by the locations of the nonzero elements of the perturbation matrix. Numerical verification of this conjecture is provided in Appendix D for the model discussed in Section $\mathbf{IV}$.
@@ IV Application in a concrete model
While many models have been proposed to realize EPs, we select a non-Hermitian supersymmetric (SUSY) array model [PhysRevA.101.033820] to illustrate our theoretical framework. This model is ideally suited for this purpose, as it allows for the construction of EPs of any order. Specially, the Hamiltonian for an $N$-site non-Hermitian SUSY array is
@@
Figure 1: Absolute eigenenergy splitting $|\Delta E|$ for the 4-site non-Hermitian SUSY array under five different perturbations $H_{p1}$ to $H_{p5}$. The blue, magenta, green, orange and red dots represent numerical results for $H_{p1}$, $H_{p2}$, $H_{p3}$, $H_{p4}$ and $H_{p5}$, respectively. Solid curves crossing the dots correspond to the analytical predictions $|\Delta E|=\epsilon^{1/4}$, $\epsilon^{1/3}$, $\epsilon^{1/2}$, $\epsilon$, and 0, respectively. Parameters are $\omega_{0}=0$, $\gamma=1$, and $J=1$.
@@ Appendix A A brief review of the Jordan chain structure
Consider an $N$th-order EP at energy $E_{0}$, described by an $N\times N$ non-Hermitian Hamiltonian $H_{N}$. This Hamiltonian possesses a unique eigenstate $|u_{0}\rangle$ and, as shown in Ref. [seyranian2003multiparameter], N1N-1 associated vectors $\{|u_{1}\rangle,\dots,|u_{N-1}\rangle\}$ that together form a Jordan chain satisfying:
@@
When $\frac{1}{2}(N-1)<j<N-1$, or equivalently Nj1<j<N1N-j-1<j<N-1, there exists an integer $t<j$ such that $s_{i}\not=\lambda^{i}$ for $i=j,j-1,\cdots,t$. This means that Eq. (32) cannot be reduced to Eq. (33). Consequently, it is difficult to solve Eq. (32) analytically. Numerical results in appendix D show that a perturbation in this regime typically induces multiple split eigenvalues, all of which scale as $\epsilon^{\frac{1}{N-j}}$.
Appendix D Perturbations beyond the four types discussed in section $\mathbf{III}$
In Section $\mathbf{III}$, we discuss four types of perturbations that can be solved analytically. For more general cases, we conjecture that the spectral splitting exponent is determined by the smallest index $j$ among the nonzero elements of the perturbation matrix, because if we consider all non-zero elements separately, those with the smallest $j$ obviously governs the spectral splitting.
Figure 2: Spectral splitting induced by (a) $\hat{H}_{p6}$ and (b) $\hat{H}_{p7}$. The orange dots represent numerical results. The three lines in (a) correspond to $\ln|\Delta E|=\frac{1}{4}\ln\epsilon+C$ with $C=0.07$, $0.008$, and 0.09-0.09, respectively. The two lines in (b) correspond to $\ln|\Delta E|=\frac{1}{2}\ln\epsilon+\frac{1}{2}\ln\frac{3\pm\sqrt{5}}{2}$. Parameters are $\omega_{0}=0$, $\gamma=1$, and $J=1$.
In this section, we numerically verify this conjecture using the non-Hermitian SUSY array model from Section $\mathbf{IV}$.
First, we investigate the case that the Hamiltonian $\hat{H}_{4}$ is perturbed by $\hat{H}_{p6}=\hat{H}_{p1}+\hat{H}_{p2}$. According to Eq. (24), the minimum index $j$ for nonzero elements in the perturbation matrix is $j_{min}=0$ . The numerical results in Fig. 2 show that the splitting of eigenenergy for all three branches scales as $\epsilon^{\frac{1}{N-j_{min}}}=\epsilon^{\frac{1}{4}}$ for small $\epsilon$, confirming the predicted exponent.
Next we examine the perturbation
$\hat{H}_{p7}$ $=$ $\frac{\epsilon J}{2\sqrt{3}}\left[2\left(c_{1}^{\dagger}c_{2}+c_{3}^{\dagger}c_{4}\right)-i\left(c_{1}^{\dagger}c_{3}-c_{2}^{\dagger}c_{4}\right)\right$. (36)
$\left.-\sqrt{3}\left(c_{1}^{\dagger}c_{4}-c_{2}^{\dagger}c_{3}\right)\right],$
with its corresponding perturbation matrix given by
$\epsilon\mathcal{H}_{1}^{p7}={\color[rgb]{0,0,0}\epsilon}\begin{pmatrix}0&0&0&0\\ J^{2}&0&0&0\\ 0&J^{2}&0&0\\ 0&0&J^{2}&0\end{pmatrix}$. (37)
[…]